# Measurement-aligned Flow for Inverse Problem

### v1

@@ 5.3 Image restoration
Inverse problems with Gaussian noise (known variance). Quantitative results for inverse problems with Gaussian noise of known variance are shown in Table 2. MAS consistently demonstrates superior performance across most tasks, notably achieving significantly lower FID scores.
@@
PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$
SR 8×8$\times$8 × DDNM 20.23 0.592 0.597 368.76 24.07 0.448 0.725 265.62
DDRM 20.26 0.592 0.602 360.79 24.11 0.451 0.729 268.48
$\Pi$GDM 20.31 0.575 0.616 338.50 24.04 0.432 0.726 222.93
RED-Diff 20.72 0.590 0.656 400.16 24.10 0.478 0.697 315.58
MAS (ours) 25.68 0.352 0.783 124.66 25.63 0.364 0.768 132.11
Deblurring DDNM 5.11 0.845 0.004 500.61 6.30 0.764 0.049 442.27
DDRM 5.11 0.845 0.004 499.67 6.30 0.764 0.049 446.33
$\Pi$GDM 5.10 0.845 0.004 497.12 6.30 0.769 0.048 451.63
RED-Diff 5.11 0.844 0.004 501.76 7.17 0.755 0.073 506.21
MAS (ours) 27.10 0.314 0.788 111.51 23.47 0.409 0.742 138.48
Inverse Problems with Non-Gaussian Noise (Unknown Strength). Quantitative evaluations for linear inverse problems with unknown, non-Gaussian noise are presented in Table 3. MAS outperforms baseline methods across all metrics, highlighting the effectiveness of our approach in handling unknown noise conditions.
@@
Inverse problems with non-differentiable measurements. MAS is also capable of solving inverse problem with non-differentiable measurements, such as JPEG restoration and quantization. Results in Table 4 and Fig. 5 show that MAS achieves state-of-the-art performance without relying on the forward operator or knowledge of the degradation strength.
Table 4: Quantitative evaluation of solving JPEG restoration and Quantization. We set $k=1.0$ for $\mathrm{QF}=5$ and $k=3.0$ for $\mathrm{QF}=2$, and $k=0.5$ for quantization. Both $\Pi$GDM and MAS use $\mathrm{NFE}=20$, which yields the best performance (among $\mathrm{NFE}=20$ and $\mathrm{NFE}=100$). Notably, our method (MAS) does not require access to the forward operator or the strength of degration.
@@
PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$
$\Pi$GDM 32.25 0.203 0.887 130.57 29.35 0.270 0.838 147.93 35.15 0.150 0.922 112.40
MAS (ours) 32.31 0.225 0.898 101.26 29.73 0.277 0.862 114.83 34.98 0.149 0.937 69.61
@@ 6 Related work
Diffusion models have also been successfully applied to linear inverse problems, including, compressed-sensing MRI (CS-MRI), and computed tomography (CT) (kadkhodaie2021stochastic, ; song2020score, ; chung2022improving, ; kawar2022denoising, ; song2021solving, ). They have also been extended to non-linear inverse problems such as Fourier phase retrieval, nonlinear deblurring, HDR, and JPEG restoration (chung2022diffusion, ; song2023pseudoinverse, ; chung2023parallel, ; mardani2023variational, ). RED-diff (mardani2023variational, ) leverages stochastic optimization to approximate the true posterior. Besides, Chung et al. (chung2023decomposed, ) combine the diffusion sampling and Krylov subspace methods, which negates the need to compute the manifold-constrained gradient. InverseBench (zheng2025inversebench, ) presents a benchmark for critical scientific applications, which present structural challenges that differ significantly from natural image restoration tasks.
In recent years, diffusion bridge models have been successfully applied to various inverse problems (liu2023i2sb, ; zhou2023denoising, ; albergo2023stochastic2, ). These models are more task-specific and directly build probabilistic paths from corrupted to clean images, showing strong performance. Exploring alignment techniques within diffusion bridges is a promising direction for further improvement.
@@ 7 Conclusion
MAS improves upon existing methods by explicitly aligning the sampling process with measurement data, offering a broader optimization perspective that generalizes approaches like DDNM and DAPS. Beyond the noise-free case, MAS can be extended to: (1) known Gaussian noise, (2) unknown or non-Gaussian noise through adaptive parameterization, and (3) non-differentiable degradations (e.g., JPEG) by decoupling the forward operator from sampling. Extensive experiments show that MAS consistently outperforms state-of-the-art methods across a wide range of inverse problems.
Limitations. (i) While MAS can handle non-differentiable measurements like JPEG restoration, it does not support general non-linear inverse problems. (ii) Tweedie’s formula boys2023tweedie could be used to improve the MAS prior by better estimating the covariance of $p(x_{0}\mid x_{t})$. Also, enhancing the likelihood model $p(y\mid\tilde{x}_{0})$ may further improve performance.
@@ Appendix A Proofs
Write
$\Sigma_{y}\;\coloneqq\;r_{t}^{2}\,HH^{\mathsf{T}}+\sigma_{y}^{2}\,\mathbb{I}_{m},\qquad\Sigma_{0}\;\coloneqq\;\rho_{t}^{2}\,\mathbb{I}_{d},\qquad\mu_{0}\;\coloneqq\;\hat{x}_{0}^{\theta}.$ (21)
Because both the likelihood $p(y\mid\tilde{x}_{0})=\mathcal{N}\!\bigl{(}H\tilde{x}_{0},\Sigma_{y}\bigr{)}$ and the prior $p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta})=\mathcal{N}\!\bigl{(}\mu_{0},\Sigma_{0}\bigr{)}$ are multivariate Gaussians, their product is proportional to another Gaussian. Using Bayes’ theorem and taking logarithms,
$\log p(\tilde{x}_{0}\mid y,\hat{x}_{0}^{\theta})=-\tfrac{1}{2}\!\Bigl{[}(y-H\tilde{x}_{0})^{\mathsf{T}}\Sigma_{y}^{-1}(y-H\tilde{x}_{0})+(\tilde{x}_{0}-\mu_{0})^{\mathsf{T}}\Sigma_{0}^{-1}(\tilde{x}_{0}-\mu_{0})\Bigr{]}+C.$ (22)
Expand the quadratic terms and collect everything that depends on $\tilde{x}_{0}$:
$\log p(\tilde{x}_{0}\mid y,\hat{x}_{0}^{\theta})=-\tfrac{1}{2}\tilde{x}_{0}^{\mathsf{T}}\bigl{(}\Sigma_{0}^{-1}+H^{\mathsf{T}}\Sigma_{y}^{-1}H\bigr{)}\tilde{x}_{0}+\tilde{x}_{0}^{\mathsf{T}}\bigl{(}\Sigma_{0}^{-1}\mu_{0}+H^{\mathsf{T}}\Sigma_{y}^{-1}y\bigr{)}+C^{\prime}.$ (23)
The exponent is a (shifted) quadratic form in $\tilde{x}_{0}$; completing the square yields
$\log p(\tilde{x}_{0}\mid y,\hat{x}_{0}^{\theta})=-\tfrac{1}{2}\!\bigl{(}\tilde{x}_{0}-\mu_{\mathrm{post}}\bigr{)}^{\mathsf{T}}\Sigma_{\mathrm{post}}^{-1}\bigl{(}\tilde{x}_{0}-\mu_{\mathrm{post}}\bigr{)}+C^{\prime\prime},$ (24)
where
$\boxed{\;\Sigma_{\mathrm{post}}=\bigl{(}\Sigma_{0}^{-1}+H^{\mathsf{T}}\Sigma_{y}^{-1}H\bigr{)}^{-1}\;}\quad\text{and}\quad\boxed{\;\mu_{\mathrm{post}}=\Sigma_{\mathrm{post}}\,\bigl{(}\Sigma_{0}^{-1}\mu_{0}+H^{\mathsf{T}}\Sigma_{y}^{-1}y\bigr{)}\;}.$ (25)
Substituting $\Sigma_{0}^{-1}=\rho_{t}^{-2}\mathbb{I}_{d}$ and $\Sigma_{y}^{-1}=(r_{t}^{2}HH^{\mathsf{T}}+\sigma_{y}^{2}\mathbb{I}_{m})^{-1}$ gives precisely the expressions in (9). Hence the posterior distribution is the multivariate Gaussian stated in the proposition. ∎
A.2 Proof of effieient linear solves in Eq. 13
@@
$W\coloneqq\eta_{1}HH^{\mathsf{T}}+\eta_{2}\mathbb{I}_{m},\qquad Y\coloneqq\mathbb{I}_{d}+H^{\mathsf{T}}W^{-1}H.$
@@
[…]
@@
$W=\eta_{1}U\Sigma\Sigma^{\mathsf{T}}U^{\mathsf{T}}+\eta_{2}U\mathbb{I}_{m}U^{\mathsf{T}}=U\!\bigl{(}\eta_{1}\Sigma\Sigma^{\mathsf{T}}+\eta_{2}\mathbb{I}_{m}\bigr{)}U^{\mathsf{T}}.$
Because $U$ is orthogonal, $W^{-1}$ is obtained by inverting the diagonal middle matrix: $(\eta_{1}\Sigma\Sigma^{\mathsf{T}}+\eta_{2}\mathbb{I}_{m})^{-1}=\operatorname{diag}\!\bigl{(}\frac{1}{\eta_{1}s_{i}^{2}+\eta_{2}}\bigr{)}_{i=1}^{m}.$ Substituting yields the first identity in (26).
@@
$Y=\mathbb{I}_{d}+H^{\mathsf{T}}W^{-1}H=V\Sigma^{\mathsf{T}}U^{\mathsf{T}}\bigl{[}U\operatorname{diag}\!\bigl{(}\tfrac{1}{\eta_{1}s_{i}^{2}+\eta_{2}}\bigr{)}U^{\mathsf{T}}\bigr{]}U\Sigma V^{\mathsf{T}}+\mathbb{I}_{d},$
and simplify with $U^{\mathsf{T}}U=\mathbb{I}_{m}$:
$Y=V\!\Bigl{[}\Sigma^{\mathsf{T}}\!\operatorname{diag}\!\bigl{(}\tfrac{1}{\eta_{1}s_{i}^{2}+\eta_{2}}\bigr{)}\!\Sigma+\mathbb{I}_{d}\Bigr{]}V^{\mathsf{T}}.$
Because $\Sigma^{\mathsf{T}}\!\operatorname{diag}\!\bigl{(}\tfrac{1}{\eta_{1}s_{i}^{2}+\eta_{2}}\bigr{)}\Sigma$ is diagonal with $i^{\text{th}}$ entry $\tfrac{s_{i}^{2}}{\eta_{1}s_{i}^{2}+\eta_{2}}$, the bracketed matrix is diagonal and hence trivial to invert, giving the second identity in (26). ∎
A.3 Proof of Eq. 14
@@
$W\;=\;\eta_{1}\,HH^{\mathsf{T}}+\eta_{2}\,\mathbb{I}_{m},\qquad Y\;=\;\mathbb{I}_{d}+H^{\mathsf{T}}W^{-1}H.$
For any $y\in\mathbb{R}^{m}$ and $\hat{x}_{0}^{\theta}\in\mathbb{R}^{d}$ consider the strictly convex quadratic
$\mathcal{L}(\tilde{x}_{0})\;=\;\bigl{\|}\tilde{x}_{0}-\hat{x}_{0}^{\theta}\bigr{\|}_{2}^{2}\;+\;\bigl{\|}y-H\tilde{x}_{0}\bigr{\|}_{W^{-1}}^{2},\qquad\|v\|_{W^{-1}}^{2}=v^{\mathsf{T}}W^{-1}v.$
@@
$\tilde{x}_{0}^{\mathrm{pm}}\;=\;Y^{-1}\!\bigl{[}\hat{x}_{0}^{\theta}+H^{\mathsf{T}}W^{-1}y\bigr{]}.$ (27)
@@
[…]
@@
$\bigl{(}\mathbb{I}_{d}+H^{\mathsf{T}}W^{-1}H\bigr{)}\tilde{x}_{0}\;=\;\hat{x}_{0}^{\theta}+H^{\mathsf{T}}W^{-1}y,\qquad\text{that is, }Y\,\tilde{x}_{0}=\hat{x}_{0}^{\theta}+H^{\mathsf{T}}W^{-1}y.$
Because $\eta_{2}>0$ implies $W\succ 0$, we have $W^{-1}\succ 0$ and hence $Y=\mathbb{I}_{d}+H^{\mathsf{T}}W^{-1}H\succ 0$; thus $Y$ is invertible and (27) follows.
Finally, the Hessian of $\mathcal{L}$ is $2Y\succ 0$, so $\mathcal{L}$ is strictly convex and the stationary point (27) is indeed its unique global minimiser. ∎
A.4 Proof of Remark 2.
@@
We start with the posterior expectation
$\mathbb{E}[\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y]=\left(\frac{1}{\rho_{t}^{2}}\mathbb{I}+\frac{1}{r_{t}^{2}}H^{\dagger}H\right)^{-1}\left(\frac{1}{\rho_{t}^{2}}\hat{x}_{0}^{\theta}+\frac{1}{r_{t}^{2}}H^{\dagger}y\right).$ (28)
We want to examine the limit as
$\frac{r_{t}}{\rho_{t}}\rightarrow 0,$ (29)
which implies that $r_{t}^{2}/\rho_{t}^{2}\rightarrow 0$. To make this explicit, set
$\epsilon=\frac{r_{t}^{2}}{\rho_{t}^{2}},$ (30)
so that $\epsilon\to 0$. Then we can write
$\frac{1}{\rho_{t}^{2}}=\frac{1}{r_{t}^{2}}\,\epsilon,$ (31)
and the expression becomes
$\mathbb{E}[\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y]=\left(\frac{\epsilon}{r_{t}^{2}}\,\mathbb{I}+\frac{1}{r_{t}^{2}}H^{\dagger}H\right)^{-1}\left(\frac{\epsilon}{r_{t}^{2}}\hat{x}_{0}^{\theta}+\frac{1}{r_{t}^{2}}H^{\dagger}y\right).$ (32)
Factor out $1/r_{t}^{2}$ from both terms:
$\mathbb{E}[\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y]=\left(\frac{1}{r_{t}^{2}}\Bigl{(}\epsilon\,\mathbb{I}+H^{\dagger}H\Bigr{)}\right)^{-1}\left(\frac{1}{r_{t}^{2}}\Bigl{(}\epsilon\,\hat{x}_{0}^{\theta}+H^{\dagger}y\Bigr{)}\right).$ (33)
The $1/r_{t}^{2}$ factors cancel (since the inverse of $aM$ is $M^{-1}/a$), yielding
$\mathbb{E}[\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y]=\left(\epsilon\,\mathbb{I}+H^{\dagger}H\right)^{-1}\left(\epsilon\,\hat{x}_{0}^{\theta}+H^{\dagger}y\right).$ (34)
To analyze the limit as $\epsilon\to 0$, decompose the space into two orthogonal components:
@@
$v=Pv+(I-P)v.$ (36)
Notice that $H^{\dagger}y$ lies in the row space (i.e. $P\,H^{\dagger}y=H^{\dagger}y$) and that $\hat{x}_{0}^{\theta}$ can be decomposed as
$\hat{x}_{0}^{\theta}=P\hat{x}_{0}^{\theta}+(I-P)\hat{x}_{0}^{\theta}.$ (37)
Since the eigenvalues of $P$ are 0 and 1, the matrix $\epsilon\,I+P$ has eigenvalues $\epsilon$ (on the nullspace of $P$) and $1+\epsilon$ (on the row space). Hence, its inverse acts as:
  • Multiplication by $1/\epsilon$ on the nullspace,
  • Multiplication by $1/(1+\epsilon)$ on the row space.
@@
$\left(\epsilon\,I+P\right)^{-1}\left(H^{\dagger}y+\epsilon\,\hat{x}_{0}^{\theta}\right)=\frac{1}{1+\epsilon}\bigl{(}H^{\dagger}y+\epsilon\,P\hat{x}_{0}^{\theta}\bigr{)}+\frac{1}{\epsilon}\bigl{(}\epsilon\,(I-P)\hat{x}_{0}^{\theta}\bigr{)}.$ (38)
@@
$\frac{1}{1+\epsilon}\,H^{\dagger}y+\frac{\epsilon}{1+\epsilon}\,P\hat{x}_{0}^{\theta}+(I-P)\hat{x}_{0}^{\theta}.$ (39)
Now, taking the limit as $\epsilon\to 0$:
  • $\frac{1}{1+\epsilon}\to 1$,
  • $\frac{\epsilon}{1+\epsilon}\to 0$.
@@
$\lim_{\epsilon\to 0}\mathbb{E}[\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y]=H^{\dagger}y+(I-P)\hat{x}_{0}^{\theta}.$ (40)
@@
$H^{\dagger}y+\hat{x}_{0}^{\theta}-H^{\dagger}H\,\hat{x}_{0}^{\theta}=\hat{x}_{0}^{\theta}+H^{\dagger}\bigl{(}y-H\hat{x}_{0}^{\theta}\bigr{)}.$ (41)
Thus, in the limit where $\frac{r_{t}}{\rho_{t}}\to 0$, we indeed have
$\mathbb{E}[\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y]=\hat{x}_{0}^{\theta}+H^{\dagger}\bigl{(}y-H\hat{x}_{0}^{\theta}\bigr{)}.$ (42)
This shows that, as the relative measurement noise $r_{t}$ becomes much smaller compared to the prior uncertainty $\rho_{t}$, the posterior expectation is the projection of $\hat{x}_{0}^{\theta}$ onto the subspace $\{x:Hx=y\}$. ∎

### v2

@@ 5.3 Image restoration
Inverse problems with Gaussian noise (known variance). Quantitative results for inverse problems with Gaussian noise of known variance are shown in Table˜2. MAS consistently demonstrates superior performance across most tasks, notably achieving significantly higher PSNRs. The table summarizes 5 tasks, 4 restoration quality metrics, and 2 datasets, resulting in a total of 40 evaluations. MAS demonstrates superior performance in 29 out of the 40 cases. Notably, MAS achieves improvements of more than 1 dB in 5 out of 10 instances.
@@
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
SR $8\times$ DDNM 13.02 0.289 0.710 377.54 18.61 0.492 0.495 268.36
DDRM 16.06 0.506 0.629 351.69 19.74 0.545 0.463 218.38
[…]
RED-Diff 14.21 0.357 0.668 342.09 19.47 0.596 0.416 224.07
MAS (ours) 20.05 0.605 0.390 129.80 20.10 0.591 0.395 137.57
Inpaint (Box) DDNM 15.55 0.248 0.533 247.99 18.60 0.621 0.341 147.80
DDRM 20.27 0.599 0.350 142.01 18.74 0.589 0.423 199.14
[…]
RED-Diff 15.75 0.287 0.523 255.85 19.13 0.638 0.338 159.99
MAS (ours) 22.78 0.723 0.244 90.15 19.13 0.581 0.407 138.56
Inverse Problems with Non-Gaussian Noise (Unknown Strength). Quantitative evaluations for linear inverse problems with unknown, non-Gaussian noise are presented in Table˜3. MAS consistently outperforms baseline methods, highlighting the effectiveness of our approach in handling unknown noise conditions.
@@
Inverse problems with non-differentiable measurements. MAS is also capable of solving inverse problems with non-differentiable measurements, such as JPEG restoration and quantization. Results in Table˜4 and Fig.˜5 show that MAS achieves state-of-the-art performance without relying on the forward operator or knowledge of the degradation strength.
Computational time analysis. The computational efficiency of MAS is comparable to DDNM and substantially higher than DAPS. For example, on the SR task using the FFHQ-256 dataset with 200 diffusion steps, the non-parallel single-image sampling time for both DDNM and MAS is only 8 seconds per image, whereas DAPS requires 67 seconds (test were conducted on the same NVIDIA A6000 GPU).
[…]
@@
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
[…]
MAS (ours) $\mathbf{26.30}$ $\mathbf{0.787}$ $\mathbf{0.281}$ $101.24$ $\mathbf{23.72}$ $\mathbf{0.772}$ $\mathbf{0.335}$ $114.85$ $28.97$ $\mathbf{0.837}$ $\mathbf{0.196}$ $\mathbf{69.61}$
@@ 6 Related work
Diffusion models have also been successfully applied to linear inverse problems, including, compressed-sensing MRI (CS-MRI), and computed tomography (CT) (kadkhodaie2021stochastic, ; song2020score, ; chung2022improving, ; kawar2022denoising, ; song2021solving, ). They have also been extended to non-linear inverse problems such as Fourier phase retrieval, nonlinear deblurring, HDR, and JPEG restoration (chung2022diffusion, ; song2023pseudoinverse, ; chung2023parallel, ; mardani2023variational, ).
Methods to solve inverse problems include linear projection methods (wang2022zero, ; kawar2022denoising, ; dou2024diffusion, ), Monte Carlo sampling (wu2023practical, ; phillips2024particle, ), variational inference (feng2023score, ; mardani2023variational, ; janati2024divide, ), optimization-based approaches (song2023solving, ; zhu2023denoising, ; li2024decoupled, ; wang2024dmplug, ; alkhouri2024sitcom, ; he2023manifold, ), and Diffusion Posterior Sampling (DPS) (zhang2024improving, ; chung2022diffusion, ; song2023loss, ; yu2023freedom, ; rout2024beyond, ; yang2024guidance, ; bansal2023universal, ; boys2023tweedie, ; song2023pseudoinverse, ; ho2022classifier, ). Besides, InverseBench (zheng2025inversebench, ) presents a benchmark for critical scientific applications, which present structural challenges that differ significantly from natural image restoration tasks.
@@ 7 Conclusion
MAS improves upon existing methods by explicitly aligning the sampling process with measurement data, offering a broader optimization perspective that generalizes approaches like DDNM and DAPS. Beyond the noise-free case, MAS can be extended to: (1) known Gaussian noise, (2) unknown or non-Gaussian noise through adaptive parameterization, and (3) non-differentiable degradations (e.g., JPEG) by decoupling the forward operator from sampling. Extensive experiments show that MAS consistently outperforms state-of-the-art methods across a wide range of inverse problems. While MAS can handle non-differentiable measurements like JPEG restoration, it does not support general non-linear inverse problems, it’s also promising to ‘calibrate’ the noise introduced into $x_{t}$, such that maximizing the consistency to measurement.
@@ Appendix A Proofs
Let $x\equiv x_{\epsilon}$. The prior and likelihood are
$p(x\mid x_{t})=\mathcal{N}(m_{0\mid t},\,C_{0\mid t}),\qquad p(y\mid x)=\mathcal{N}(Hx,\,R),$
with $R=\sigma_{y}^{2}I_{m}+\sigma_{\epsilon}^{2}HH^{\top}$. Denote $m\coloneqq m_{0\mid t}$ and $C\coloneqq C_{0\mid t}$.
The posterior is, up to normalization,
$p(x\mid x_{t},y)\;\propto\;\exp\!\left(-\tfrac{1}{2}(x-m)^{\top}C^{-1}(x-m)-\tfrac{1}{2}(y-Hx)^{\top}R^{-1}(y-Hx)\right).$
Expanding the exponent and collecting terms in $x$ gives
$-\tfrac{1}{2}\Big[x^{\top}C^{-1}x-2x^{\top}C^{-1}m+m^{\top}C^{-1}m+x^{\top}H^{\top}R^{-1}Hx-2x^{\top}H^{\top}R^{-1}y+y^{\top}R^{-1}y\Big]$
$=-\tfrac{1}{2}\Big[x^{\top}(C^{-1}+H^{\top}R^{-1}H)x-2x^{\top}(C^{-1}m+H^{\top}R^{-1}y)\Big]+\text{(terms independent of $x$)}.$
This is the quadratic form of a Gaussian in $x$ with precision
$\Lambda\;=\;C^{-1}+H^{\top}R^{-1}H,$
and natural parameter
$\eta\;=\;C^{-1}m+H^{\top}R^{-1}y.$
Therefore the posterior is Gaussian $\mathcal{N}(\mu_{\text{post}},\,\Sigma_{\text{post}})$ with
$\Sigma_{\text{post}}=\Lambda^{-1}=\big(C^{-1}+H^{\top}R^{-1}H\big)^{-1},\qquad\mu_{\text{post}}=\Sigma_{\text{post}}\,\eta=\big(C^{-1}+H^{\top}R^{-1}H\big)^{-1}\big(C^{-1}m+H^{\top}R^{-1}y\big).$
Restoring the original notation gives equation 9. ∎
A.2 Proof of effieient linear solves in Eq.˜38
@@
$W\coloneqq\eta_{1}HH^{\mathsf{T}}+\eta_{2}\mathbb{I},\qquad Y\coloneqq\mathbb{I}+H^{\mathsf{T}}W^{-1}H.$
@@
$W^{-1}=U\,\operatorname{diag}\!\Bigl(\tfrac{1}{\eta_{1}s_{i}^{2}+\eta_{2}}\Bigr)_{i=1}^{m}U^{\mathsf{T}},\qquad Y^{-1}=V\,\operatorname{diag}\!\Bigl(\tfrac{1}{\,1+s_{i}^{2}/(\eta_{1}s_{i}^{2}+\eta_{2})\,}\Bigr)_{i=1}^{d}V^{\mathsf{T}}.$ (14)
@@
$W=\eta_{1}U\Sigma\Sigma^{\mathsf{T}}U^{\mathsf{T}}+\eta_{2}U\mathbb{I}U^{\mathsf{T}}=U\!\bigl(\eta_{1}\Sigma\Sigma^{\mathsf{T}}+\eta_{2}\mathbb{I}\bigr)U^{\mathsf{T}}.$
Because $U$ is orthogonal, $W^{-1}$ is obtained by inverting the diagonal middle matrix: $(\eta_{1}\Sigma\Sigma^{\mathsf{T}}+\eta_{2}\mathbb{I})^{-1}=\operatorname{diag}\!\bigl(\frac{1}{\eta_{1}s_{i}^{2}+\eta_{2}}\bigr)_{i=1}^{m}.$ Substituting yields the first identity in equation 14.
@@
$Y=\mathbb{I}+H^{\mathsf{T}}W^{-1}H=V\Sigma^{\mathsf{T}}U^{\mathsf{T}}\bigl[U\operatorname{diag}\!\bigl(\tfrac{1}{\eta_{1}s_{i}^{2}+\eta_{2}}\bigr)U^{\mathsf{T}}\bigr]U\Sigma V^{\mathsf{T}}+\mathbb{I},$
and simplify with $U^{\mathsf{T}}U=\mathbb{I}$:
$Y=V\!\Bigl[\Sigma^{\mathsf{T}}\!\operatorname{diag}\!\bigl(\tfrac{1}{\eta_{1}s_{i}^{2}+\eta_{2}}\bigr)\!\Sigma+\mathbb{I}\Bigr]V^{\mathsf{T}}.$
Because $\Sigma^{\mathsf{T}}\!\operatorname{diag}\!\bigl(\tfrac{1}{\eta_{1}s_{i}^{2}+\eta_{2}}\bigr)\Sigma$ is diagonal with $i^{\text{th}}$ entry $\tfrac{s_{i}^{2}}{\eta_{1}s_{i}^{2}+\eta_{2}}$, the bracketed matrix is diagonal and hence trivial to invert, giving the second identity in equation 14. ∎
A.3 Proof of Eq.˜6
@@
$W\;=\;\eta_{1}\,HH^{\mathsf{T}}+\eta_{2}\,\mathbb{I},\qquad Y\;=\;\mathbb{I}+H^{\mathsf{T}}W^{-1}H.$
For any $y\in\mathbb{R}^{m}$ and $m_{0\mid t}\in\mathbb{R}^{d}$ consider the strictly convex quadratic
$\mathcal{L}(x_{0})\;=\;\bigl\|x_{0}-m_{0\mid t}\bigr\|_{2}^{2}\;+\;\bigl\|y-H\tilde{x}_{0}\bigr\|_{W^{-1}}^{2},\qquad\|v\|_{W^{-1}}^{2}=v^{\mathsf{T}}W^{-1}v.$
@@
$\tilde{x}_{0}^{*}\;=\;Y^{-1}\!\bigl[m_{0\mid t}+H^{\mathsf{T}}W^{-1}y\bigr].$ (15)
@@
$\nabla_{\tilde{x}_{0}}\mathcal{L}\;=\;2\bigl(\tilde{x}_{0}-m_{0\mid t}\bigr)\;-\;2H^{\mathsf{T}}W^{-1}\!\bigl(y-H\tilde{x}_{0}\bigr).$
@@
$\bigl(\mathbb{I}+H^{\mathsf{T}}W^{-1}H\bigr)\tilde{x}_{0}\;=\;m_{0\mid t}+H^{\mathsf{T}}W^{-1}y,\qquad\text{that is, }Y\,\tilde{x}_{0}=m_{0\mid t}+H^{\mathsf{T}}W^{-1}y.$
Because $\eta_{2}>0$ implies $W\succ 0$, we have $W^{-1}\succ 0$ and hence $Y=\mathbb{I}+H^{\mathsf{T}}W^{-1}H\succ 0$; thus $Y$ is invertible and equation 15 follows.
Finally, the Hessian of $\mathcal{L}$ is $2Y\succ 0$, so $\mathcal{L}$ is strictly convex and the stationary point equation 15 is indeed its unique global minimiser. ∎
A.4 Proof of Remark˜1.
@@
As $\eta_{2}=0$,
$x_{0}^{*}=\left(\eta\,\mathbb{I}+H^{\dagger}H\right)^{-1}\left(\eta_{1}\,m_{0\mid t}+H^{\dagger}y\right).$ (16)
To analyze the limit as $\eta_{1}\to 0$, decompose the space into two orthogonal components:
@@
v=Pv+(IP)v.v=Pv+(I-P)v. (18)
Notice that $H^{\dagger}y$ lies in the row space (i.e. $P\,H^{\dagger}y=H^{\dagger}y$) and that $m_{0\mid t}$ can be decomposed as
$m_{0\mid t}=Pm_{0\mid t}+(I-P)m_{0\mid t}.$ (19)
Since the eigenvalues of $P$ are 0 and 1, the matrix $\eta_{1}\,I+P$ has eigenvalues $\eta_{1}$ (on the nullspace of $P$) and $1+\eta_{1}$ (on the row space). Hence, its inverse acts as:
  • Multiplication by $1/\eta_{1}$ on the nullspace,
  • Multiplication by $1/(1+\eta_{1})$ on the row space.
@@
$\left(\eta_{1}\,I+P\right)^{-1}\left(H^{\dagger}y+\eta_{1}\,m_{0\mid t}\right)=\frac{1}{1+\epsilon}\bigl(H^{\dagger}y+\eta_{1}\,Pm_{0\mid t}\bigr)+\frac{1}{\eta_{1}}\bigl(\eta_{1}\,(I-P)m_{0\mid t}\bigr).$ (20)
@@
$\frac{1}{1+\eta_{1}}\,H^{\dagger}y+\frac{\eta_{1}}{1+\eta_{1}}\,Pm_{0\mid t}+(I-P)m_{0\mid t}.$ (21)
Now, taking the limit as $\eta_{1}\to 0$:
  • $\frac{1}{1+\eta_{1}}\to 1$,
  • $\frac{\eta_{1}}{1+\eta_{1}}\to 0$.
@@
$\lim_{\eta_{1}\to 0}x_{0}^{*}=H^{\dagger}y+(I-P)m_{0\mid t}.$ (22)
@@
$H^{\dagger}y+m_{0\mid t}-H^{\dagger}H\,m_{0\mid t}=m_{0\mid t}+H^{\dagger}\bigl(y-Hm_{0\mid t}\bigr).$ (23)
Thus, in the limit where $\eta_{1}\to 0$, we indeed have
$x_{0}^{*}=m_{0\mid t}+H^{\dagger}\bigl(y-Hm_{0\mid t}\bigr).$ (24)
This shows that, as the relative measurement noise $\epsilon$ becomes much smaller compared to the prior uncertainty $r_{t}$, the posterior expectation is the projection of $\hat{x}_{0}^{\theta}$ onto the subspace $\{x:Hx=y\}$. ∎